Describe how will a bank creates bank money when initial demand deposit is $2000, required ratio is 5% and lending the excess reserve to three borrowers.
Answer #1

Bank creates money based on the reserve requirement. Under the fractional reserve system, bank can create money:

First Borrower:

First Bank Balance Sheet:

Assets Liabilities

Reserve:100

Loan: 1900

Deposit: 2000

Second Bank Balance Sheet

Assets Liabilities

Reserve:95

Loan: 1805

Deposit: 1900

Third Bank Balance Sheet:

Assets Liabilities

Reserve:90.25

Loan: 1714.5

Deposit: 1805

Money Multiplier = 1/rr

= 1/0.05

= 20

Money Supply = 20*2000

= $40,000